OpenStudy (anonymous):

15(2 - x) = 13(3 - x) is it -9/2

OpenStudy (anonymous):

awesome again

OpenStudy (anonymous):

30 - 15x = 39 - 13x 2x = -9 x = -9/2

OpenStudy (anonymous):

tick